# Cycles of Length 4 or 8 in Graphs with Diameter 2 and Minimum Degree at Least 3

### v3

@@ Introduction
Subcase 2A: $v_{8}=v_{1}$. Then $v_{1}v_{6}\in E(G)$. Since $\delta(G)\geq 3$, choose
$w\in N(v_{6})\setminus\{v_{1},v_{2}\}.$
By Claim 2.1, $w\notin\{v_{3},v_{4}\}$; moreover $w\neq v_{5}$ or $v_{7}$, else a $4$-cycle occurs. The vertices $v_{4}$ and $w$ are nonadjacent, or else $v_{4}-w-v_{6}-v_{1}-v_{4}$ is a $4$-cycle. Thus $\operatorname{diam}(G)=2$ yields
$z\in N(v_{4})\cap N(w),$
with $z$ distinct from all previously defined vertices. Then
$v_{7}-v_{3}-v_{1}-v_{4}-z-w-v_{6}-v_{2}-v_{5}-v_{7}$
contains a cycle of length $8$.
Subcase 2B: $v_{8}=v_{2}$. This case is handled analogously to $v_{8}=v_{1}$: using $\delta(G)\geq 3$, choose $w\in N(v_{4})\setminus\{v_{1},v_{2}\}$, giving $z\in N(w)\cap N(v_{6})$ via $\operatorname{diam}(G)=2$. Thus, obtaining an $8$-cycle.
@@
which has length $8$.
Thus, in all cases, $G$ contains a cycle of length $4$ or $8$ (see Figure 8).

### v4

@@ Introduction
Subcase 2A: $v_{8}=v_{1}$. Then $v_{1}v_{6}\in E(G)$. Since $\operatorname{diam}(G)=2$ and $v_{6}$ and $v_{7}$ are nonadjacent, we have
$N(v_{6})\cap N(v_{7})\neq\varnothing.$
Choose $x\in N(v_{6})\cap N(v_{7})$. If $x\in V(G^{\prime})\setminus\{v_{6}\}$ then a 44–cycle appears by at least one of the following:
$x=v_{1}$ $\Rightarrow v_{1}-v_{7}-v_{5}-v_{2}-v_{1},$
$x=v_{2}$ $\Rightarrow v_{2}-v_{1}-v_{3}-v_{7}-v_{2},$
$x=v_{3}$ $\Rightarrow v_{3}-v_{1}-v_{2}-v_{6}-v_{3},$
$x=v_{4}$ $\Rightarrow v_{4}-v_{1}-v_{2}-v_{6}-v_{4},$
$x=v_{5}$ $\Rightarrow v_{5}-v_{2}-v_{1}-v_{6}-v_{5}.$
Thus, $x\notin V(G^{\prime})\setminus\{v_{6}\}$ (see Figure 8).
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$x$
Figure 8: Common neighbor $x\in N(v_{6})\cap N(v_{7})$ with $x\notin V(G^{\prime})\setminus\{v_{6}\}$.
Since $v_{4}$ and $v_{5}$ are nonadjacent (by Claim 2.1),
$N(v_{4})\cap N(v_{5})\neq\varnothing,$
by the $\operatorname{diam}(G)=2$ condition.
Suppose $y\in N(v_{4})\cap N(v_{5})$. By Claim 2.1, $y\notin\{v_{3},v_{6}\}$.
Suppose $y\in\{v_{1},v_{2},v_{7},x\}$.
Thus, a 44–cycle occurs by the following:
$y=v_{1}$ $\Rightarrow v_{1}-v_{6}-v_{2}-v_{5}-v_{1},$
$y=v_{2}$ $\Rightarrow v_{2}-v_{6}-v_{1}-v_{4}-v_{2},$
$y=v_{7}$ $\Rightarrow v_{7}-v_{3}-v_{1}-v_{4}-v_{7},$
$y=x$ $\Rightarrow x-v_{5}-v_{2}-v_{6}-x.$
Therefore, $y\in N(v_{4})\cap N(v_{5})$ such that $y\notin\{v_{1},v_{2},v_{7},x\}$ (see Figure 9) .
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$xxyy
Figure 9: Common neighbor $y\in N(v_{4})\cap N(v_{5})$ with $y\notin\{v_{1},v_{2},v_{5}\}$.
But
$x-v_{7}-v_{5}-y-v_{4}-v_{1}-v_{2}-v_{6}-x$
forms a cycle of length $8$. $\diamond$
Subcase 2B: $v_{8}=v_{2}$. This case is handled analogously to $v_{8}=v_{1}$; using diameter and degree constraints, start with $v_{2}v_{4}\in E(G)$, force a $x\in N(v_{4})\cap N(v_{7})$ and $y\in N(v_{3})\cap N(v_{6})$, obtaining an 88–cycle.
@@
which has length $8$ (see Figure 10).
Thus, in all cases, $G$ contains a cycle of length $4$ or $8$.